OpenStudy (anonymous):

Factor using the Difference of Squares, if possible: 81n²-49 =

OpenStudy (dape):

The difference of squares simply says \[ (x+y)(x-y)=x^2-y^2 \] so you have to find a suitable x and y for that formula. Now \[ 81n^2=(9n)^2 \] and \[ 49=7^2 \] so \[ 81n^2-49=(9n+7)(9n-7). \]